Solve P = (r + 5t) for r
Answer by jim_thompson5910(35256)

P = (r + 5t)

P = r + 5t

P - 5t = r

r = P - 5t

So the answer is r = P - 5t

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------